OpenStudy (anonymous):

FAN AND MEDAL Solve the following system of equations: 4x - 2y - z = -5 x - 3y + 2z = 3 3x + y - 2z = -5 (0, 1, 3) (0, -1, 3) (0, 1, -3) (0, -1, -3)

OpenStudy (texaschic101):

4x - 2y - z = -5 --->(2) x - 3y + 2z = 3 --------------- 8x - 4y - 2z = -10 (result of multiplying by 2) x - 3y + 2z = 3 ---------------add 9x - 7y = - 7 x - 3y + 2z = 3 3x + y - 2z = -5 ---------------add 4x - 2y = -2 9x - 7y = -7 --->(2) 4x - 2y = -2 --->(-7) ---------------- 18x - 14y = -14 (result of multiplying by 2) -28x + 14y = 14 (result of multiplying by -7) ----------------add -10x = 0 x = 0 4x - 2y = -2 4(0) - 2y = -2 -2y = -2 y = 1 4x - 2y - z = -5 4(0) - 2(1) - z = -5 -2 - z = -5 -2 + 5 = z 3 = z check.. x - 3y + 2z = 3 0 - 3(1) + 2(3) = 3 -3 + 6 = 3 3 = 3 (correct) answer : (0,1,3)

OpenStudy (abmon98):

x=3+3y-2z 4(3+3y-2z)-2y-z=-5 12+12y-8z-2y-z=-5 12+10y-9z=-5 10y-9z=-17 3(3+3y-2z)+y-2z=-5 9+9y-6z+y-2z=-5 10y-8z=-14 -9z--8z=-17--14 -z=-3 z=3 10y-8(3)=-14 10y-24=-14 10y=-14+24 y=1 x=3+3(1)-2(3) x=3+3-6 x=0

OpenStudy (anonymous):

guess they got it :P lol